# Properties of Immersions for Systems with Multiple Limit Sets with Implications to Learning Koopman Embeddings

### v1

@@ Abstract
Linear immersions (or Koopman eigenmappings) of a nonlinear system have wide applications in prediction and control. In this work, we study the non-existence of one-to-one linear immersions for nonlinear systems with multiple omega-limit sets. While previous research has indicated the possibility of discontinuous one-to-one linear immersions for such systems, it remained uncertain whether continuous one-to-one linear immersions are attainable. Under mild conditions, we prove that any continuous one-to-one immersion to a class of systems including linear systems cannot distinguish different omega-limit sets, and thus cannot be one-to-one. Furthermore, we show that this property is also shared by approximate linear immersions learned from data as sample size increases and sampling interval decreases. Multiple examples are studied to illustrate our results.
@@
Nonlinear Systems, Immersion, Koopman Operator, Equivalent Systems
@@ 1 Introduction
Applied Koopman operator theory has drawn much attention in recent years due to its potential in the analysis, prediction, and control of nonlinear systems. The main idea behind this is fairly straightforward: As initially shown by Koopman (1931), a nonlinear system can be equivalently represented by an infinite-dimensional linear system whose states consist of single-valued observation functions of the nonlinear system. If one can further find an invariant subspace of this infinite-dimensional linear system, a finite linear representation of the nonlinear system, called the Koopman representation, can be extracted from a basis of this invariant subspace. This makes the prediction and control for the nonlinear system much easier since existing theoretical and algorithmic tools established for linear systems can now be applied to the nonlinear systems via their finite linear representations. Compared with local linearization by Taylor expansion, the Koopman representation can capture the global behaviors of the system (Mauroy et al. (2020); Brunton et al. (2021)) and thus opens up exciting possibilities in various applications, such as model reduction and control of PDEs (Kutz et al. (2018); Peitz and Klus (2020)), prediction of chaotic systems (Brunton et al. (2017)), modeling and control of soft robots (Bruder et al. (2020)), and model predictive control of nonlinear systems (Korda and Mezić (2018)).
@@
The idea behind Koopman representations and embeddings of nonlinear systems in linear (or bilinear, when there are controls) systems has been a recurring theme in the control literature, albeit under different names. Finite-dimensional embeddings correspond to finite-dimensional spaces of observables (Wang and Sontag (1989)). The Koopman representation can be interpreted as the “dual system” used in linear theory (Kalman duality) and more generally as the foundation of the duality between observability of a nonlinear system and controllability of a (generally infinite dimensional) system of observables, the adjoint system. See for example the work in Sontag and Rouchaleau (1976); Sontag (1979, 1995) on algebraic observability (strong reachability of the adjoint system, and surjective comorphisms into cosystems in the first reference) and a brief mention in Exercise 6.2.10 in the textbook Sontag (1998). A very closely related concept, but for infinite-dimensional linear systems, is “topological observability”, which amounts to the exact reachability of a dual system (Yamamoto (1981)).
@@
The primary challenge in applying Koopman operator theory to prediction and control lies in identifying a suitable Koopman representation. This involves finding a nonlinear transformation of the system states such that the transformed states evolve like a linear system. We call such a transformation a linear immersion. As a trivial example, any constant function is a linear immersion for an arbitrary system, but this linear immersion is useless in practice since it does not include any information about the original nonlinear system. Ideally, we want to find invertible linear immersions, ensuring that the trajectories of the original nonlinear system are fully characterized by its linear representation. In instances where invertible linear immersions cannot be manually derived, especially for higher-dimensional systems, numerical approximation becomes necessary. Various numerical methods have been developed to approximate linear immersions from data (Tu et al. (2014); Williams et al. (2015); Brunton et al. (2016b)). A crucial guideline for achieving low approximation error in these methods is to carefully select a domain of interest where the linear immersions are intended to be learned. In practice, for systems with multiple equilibria, a commonly mentioned insight is that a continuous one-to-one linear immersion across multiple isolated equilibria does not exist, supported by multiple analytical and numerical examples in the literature (Mauroy and Mezić (2013); Williams et al. (2015); Brunton et al. (2016a); Page and Kerswell (2019); Bakker et al. (2019, 2020)). Consequently, numerical methods are recommended to focus on learning local linear immersions within the domain of attraction of each equilibrium point. However, a recent work from Arathoon and Kvalheim (2023) challenges this insight. In particular, Arathoon and Kvalheim (2023) construct a smooth system with multiple isolated equilibria that admits a smooth one-to-one linear immersion. These positive and negative examples suggest that the non-existence of continuous one-to-one linear immersions is not solely determined by the presence of multiple isolated equilibria. To provide more accurate guidance on approximating linear immersions, it is imperative to reassess the aforementioned insight and identify the actual factors that determine the existence or non-existence of continuous one-to-one linear immersions.
@@
To address these inquiries, in this work we study the properties of continuous linear immersions for systems with multiple limit sets, and their implications on learning algorithms that approximate linear immersions from data. In particular, our contributions include:
@@
  • For systems with multiple $\omega$-limit sets, we prove that, under mild conditions, any continuous linear immersion collapses all the $\omega$-limit sets into one and thus can not be one-to-one. We then demonstrate the applicability of our results with multiple examples from the literature (Section 3).
@@
  • For the same class of systems, we show that approximate linear immersions learned with data converge to functions that are not one-to-one, as sampling time decreases and sample size increases (Section 4).
@@
A preliminary version of this work was presented at the IFAC World Congress (Liu et al. (2023)), focusing exclusively on one-to-one immersions. In this work, we extend the results in Liu et al. (2023) to encompass immersions that are not necessarily one-to-one in Section 3. Additionally, we introduce entirely new results in Sections 4 and 5.
@@
Related work: Since the presence of multiple isolated equilibria is one of the key features that distinguish nonlinear systems from linear ones, there are many discussions in the literature on the possibility of immersing a system with more than one isolated equilibria in a linear system. It is initially observed in a numerical example from Mauroy and Mezić (2013) that the approximate linear immersion over a domain that contains two equilibria becomes singular at one of the equilibria. Motivated by this observation, Williams et al. (2015) suggest that linear immersions should be approximated within the domain of attraction of each equilibrium to avoid singularities. This suggestion is supported by the work from Page and Kerswell (2019), who study a one-dimensional system with three equilibria where all the linear immersions can be derived manually. For this specific system, all the derived linear immersions become singular at one of the equilibria. Potentially motivated by these negative examples, Brunton et al. (2016a) claim that it is impossible to find one-to-one linear immersions for systems with multiple isolated equilibria. However, this claim is disproved by Bakker et al. (2019), who present a one-dimensional system with three equilibria that admits a discontinuous one-to-one linear immersion. Bakker et al. (2019) further conjecture that a linear immersion may exist but become discontinuous at the boundaries of the basins of attraction. This conjecture is again disproved by Arathoon and Kvalheim (2023), who construct a smooth system with two isolated equilibria that admits a smooth one-to-one linear immersion. Contrary to these varying claims, our work rigorously proves that continuous linear immersions cannot be one-to-one when the system has multiple isolated equilibria and satisfies specific conditions. Our results confirm that the negative examples from the literature do not admit continuous one-to-one linear immersions, while the counter example in Arathoon and Kvalheim (2023) is the only one not meeting our extra conditions and thus allowing a continuous one-to-one linear immersion. Notably, Kvalheim and Arathoon (2023) provide necessary and sufficient conditions for the existence of one-to-one linear immersions for a certain class of nonlinear systems. However, their results, when applied to systems with multiple isolated limit sets, are more restrictive than ours. This is because they assume that the system is defined over a compact invariant subset of a smooth manifold (e.g., a simple linear system like $\dot{x}=-x$ does not satisfy this assumption). Interestingly, under this assumption, our results directly imply that any systems with multiple isolated equilibria do not admit a continuous one-to-one linear immersion. This implication can not be easily discerned from the necessary and sufficient conditions presented in Kvalheim and Arathoon (2023).
@@
Notation: We denote the closure of a set $X$ by $\text{cl}(X)$. The symbols $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, and $\mathbb{R}_{>0}$ denote the real line, the set of non-negative real numbers and the set of positive real numbers. The symbols ${\mathbb{Z}}$ and $\mathbb{N}$ denote the set of integers and the set of non-negative integers.
@@ 2.1 Problem Statement
Given an initial state $\xi$ of the system in (1), we denote the $\omega$-limit set of $\xi$ in ${\mathcal{X}}$ by $\omega^{+}(\xi)$, that is the set of all $x\in{\mathcal{X}}$ satisfying that there exists a sequence $t_{n}\rightarrow\infty$ such that ${\lim_{n\to\infty}\varphi(t_{n},\xi)=x}$ (Hirsch et al. (2012)).
Next, we introduce the definition of immersion, which generalizes the notion of Koopman eigenfunctions (Mauroy et al. (2020)).
Definition 2.2
@@
where $\psi(t,F(\xi))$ is the solution of $\dot{z}=g(z)$. If the system $\dot{z}=g(z)$ above is linear, the mapping $F$ is called a linear immersion.
Remark 2.3
@@
Remark 2.4
Linear immersions are tightly related to Koopman operator theory (Brunton et al. (2016a)): A Koopman eigenfunction $F$ is a (not necessarily continuous) linear immersion that immerses $\dot{x}=f(x)$ in a one-dimensional system $\dot{z}=\lambda z$ for some $\lambda\in\mathbb{R}$. The span of the entries of any linear immersion $F$ is a Koopman invariant subspace.
Remark 2.5
@@
Remark 2.6
The term “immersion” is also widely used in the study of differentiable manifolds (see for example Lee (2012)), which is unrelated to the immersion of dynamical systems considered in this work.
@@
with state $x=(x_{1},x_{2})$. The system has an unstable equilibrium at the origin and one stable limit cycle on the unit circle, as shown by the phase portrait in Fig. 1. Let ${\mathcal{X}}=\mathbb{R}^{2}\backslash\{0\}$. Intuitively, one may think a linear immersion does not exist for this system over ${\mathcal{X}}$ since linear systems cannot have a limit cycle. However, this sytem does admit a one-to-one linear immersion over ${\mathcal{X}}$. Let $F:{\mathcal{X}}\rightarrow\mathbb{R}^{3}$ be
@@
This paper focuses on another well-known topological difference between linear and nonlinear systems, namely that a nonlinear system can have multiple isolated $\omega$-limit sets, but a linear system cannot. We wonder how the properties of linear immersions, such as injectivity, are influenced by the presence of multiple isolated $\omega$-limit sets. Many existing works (Mauroy and Mezić (2013); Brunton et al. (2016a); Page and Kerswell (2019)) claim or observe that a continuous one-to-one linear immersion does not exist when the system possesses multiple isolated equilibria, a specific type of $\omega$-limit sets. However, a formal analysis of this phenomenon is missing in the literature. The following discussion explains why it is nontrivial to prove this claim: Suppose that an immersion $F$ that maps a system $\dot{x}=f(x)$ with multiple equilibria to a linear system $\dot{z}=Az$ exists. According to (3), a one-to-one $F$ must map equilibria of $\dot{x}=f(x)$ to equilibria of the linear system $\dot{z}=Az$, and map non-equilibrium points to non-equilibrium points. However, recall that a linear system can only have one isolated equilibrium or a subspace of equilibria. If $\dot{z}=Az$ is the former, then $F$ maps all equilibria of $\dot{x}=f(x)$ to the unique equilibrium of $\dot{z}=Az$ and thus can not be one-to-one. Thus, the main challenge of proving or disproving this claim is to show if it is possible to have a one-to-one $F$ that maps $\dot{x}=f(x)$ to an immersing system $\dot{z}=Az$ with a subspace of equilibria. In this case, the only possibility is that the graph of $F$ intersects with the null space of $A$ at exactly $M$ points, with $M$ the number of equilibria of $\dot{x}=f(x)$, as demonstrated by Fig. 2. The following example shows that this is indeed possible if we allow the immersion to be discontinuous.
Example 2.8
Consider a one-dimensional system $\dot{x}=f(x)$ with $M$ isolated equilibria $\{x_{e,i}\}_{i=1}^{M}$, where $x_{e,1}<x_{e,2}<\cdots<x_{e,M}$. Assume that the solution $\varphi(t,\xi)$ is defined for any $\xi$ and $t\in\mathbb{R}$ (and thus has no finite-time blow-up both forward and backward in time). Let ${\mathcal{X}}=\mathbb{R}$. This system is immersed to the following two-dimensional linear system
@@
$\small F(x)=\begin{cases}(x,0)&f(x)=0,\\ (\varphi^{-1}(x,x_{e,1}-1),1)&x<x_{e,1},\\ \left((i+1)\varphi^{-1}\left(x,\frac{x_{e,i}+x_{e,i+1}}{2}\right),i+1\right)&x\in(x_{e,i},x_{e,i+1}),\\ ((M+1)\varphi^{-1}(x,x_{e,M}+1),M+1)&x>x_{e,M},\\ \end{cases}$ (14)
where for any $x$ and $\xi\in(x_{e,i},x_{e,i+1})$, the inverse function $\varphi^{-1}(x,\xi)$ is the time instance $t$ such that $\varphi(t,\xi)=x$. Intuitively, the equilibria of $\dot{x}=f(x)$ cut the real line into $M+1$ intervals, and the function $F$ in (14) maps all these intervals to horizontal lines in the lifted space.
To see how this immersion $F$ works in practice, we take a concrete example of $\dot{x}=f(x)$ with $M=3$
@@
The equilibria of this system include ±$\pm 1$ and 00. According to (14), a one-to-one linear immersion $F$ for this system is
$\small F(x)=\begin{cases}(x,0)&f(x)=0,\\ \left(\ln\bigg{|}\frac{3x}{-2+2x^{2}}\bigg{|},1\right)&x<-1,\\ \left(2\ln\bigg{|}\frac{3x}{-2+2x^{2}}\bigg{|},2\right)&-1<x<0,\\ \left(3\ln\bigg{|}\frac{3x}{-2+2x^{2}}\bigg{|},3\right)&0<x<1,\\ \left(4\ln\bigg{|}\frac{3x}{-2+2x^{2}}\bigg{|},4\right)&x>1.\\ \end{cases}$ (16)
Let us examine the correctness of $F$ for $x_{0}\in(0,1)$. For any $x_{0}\in(0,1)$, the solution $\varphi(t,x_{0})$ for $\dot{x}=f(x)$ in (15) is
@@
$F(\varphi(t,x_{0}))$ $=\left(3\ln\left(\frac{3x_{0}}{2x_{0}^{2}-2}\right)+3t,3\right).$ (18)
@@
Note that the graph of $F$ in (14) intersects with the subspace ×$\mathbb{R}\times\{0\}$ of equilibria of (13) at precisely $M$ points, thanks to the discontinuity. This phenomenon is also observed in continuous one-to-one linear immersions, exemplified in Arathoon and Kvalheim (2023). This work aims to elucidate the relation between the properties of linear immersions, such as injectivity and continuity, and the occurrence of multiple isolated $\omega$-limit sets, as indicated by these examples and the literature. Problem: Identify the properties of linear immersions for systems with multiple isolated $\omega$-limit sets.

### v5

@@ Abstract
Linear immersions (such as Koopman eigenfunctions) of a nonlinear system have wide applications in prediction and control. In this work, we study the properties of linear immersions for nonlinear systems with multiple omega-limit sets. While previous research has indicated the possibility of discontinuous one-to-one linear immersions for such systems, it has been unclear whether continuous one-to-one linear immersions are attainable. Under mild conditions, we prove that any continuous immersion to a class of systems including finite-dimensional linear systems collapses all the omega-limit sets, and thus cannot be one-to-one. Furthermore, we show that this property is also shared by approximate linear immersions learned from data as sample size increases and sampling interval decreases. Multiple examples are studied to illustrate our results.
@@
Nonlinear Dynamics, Immersion, Koopman Embeddings
@@ 1 Introduction
Applied Koopman operator theory has drawn much attention in recent years due to its potential in the analysis, prediction, and control of nonlinear systems. The main idea behind this is fairly straightforward: As initially shown by [13], a nonlinear system can be equivalently represented by an infinite-dimensional linear system whose states consist of observables of the nonlinear system. If a finite-dimensional invariant subspace of this infinite-dimensional linear system is found, a finite-dimensional linear representation of the nonlinear system, called the Koopman representation, can be extracted from its basis. This makes prediction and control for nonlinear systems much easier since existing theoretical and algorithmic tools established for linear systems can now be applied to nonlinear systems via their finite-dimensional linear representations. Compared with local linearization by Taylor expansion, the Koopman representation can capture global behaviors of the system [23, 8] and thus opens up exciting possibilities in various applications, such as model reduction and control of PDEs [15, 25], prediction of chaotic systems [6], modeling and control of soft robots [5], and model predictive control of nonlinear systems [14].
@@
The idea behind Koopman representations and embeddings of nonlinear systems in linear (or bilinear, when there are controls) systems has been a recurring theme in the control literature, albeit under different names. Finite-dimensional embeddings correspond to finite-dimensional spaces of observables [34]. The Koopman representation can be interpreted as the “dual system” used in linear theory (Kalman duality) and more generally as the foundation of the duality between observability of a nonlinear system and controllability of a (generally infinite-dimensional) system of observables, the adjoint system. See for example the work in [31, 27, 28] on algebraic observability (strong reachability of the adjoint system, and surjective comorphisms into cosystems in the first reference) and a brief mention in Exercise 6.2.10 in the textbook [30]. A very closely related concept, but for infinite-dimensional linear systems, is “topological observability”, which amounts to the exact reachability of a dual system [36].
@@
The primary challenge in applying Koopman operator theory to prediction and control lies in identifying a suitable Koopman representation. This involves finding a nonlinear transformation of the system states such that the transformed states evolve like a linear system. We call such a transformation a linear immersion. Scalar-valued linear immersions are also known as Koopman eigenfunctions in the literature. As a trivial example, any constant function is a linear immersion for an arbitrary system, but this linear immersion is useless in practice since it does not include any information about the original nonlinear system. Ideally, we want to find invertible linear immersions, ensuring that the trajectories of the original nonlinear system are fully characterized by its linear representation. In instances where invertible linear immersions cannot be manually derived, especially for higher-dimensional systems, numerical approximation becomes necessary. Various numerical methods have been developed to approximate linear immersions from data [9, 32, 35]. A crucial guideline for achieving low approximation error in these methods is to carefully select a domain of interest where linear immersions are intended to be learned. In practice, for systems with multiple equilibria, a commonly mentioned insight is that a continuous one-to-one linear immersion across multiple isolated equilibria does not exist, supported by multiple analytical and numerical examples in the literature [3, 4, 7, 22, 24, 35]. Consequently, numerical methods are recommended to focus on learning local linear immersions within the domain of attraction of each equilibrium. However, recent work from [2] challenges this insight. In particular, Arathoon and Kvalheim [2] construct a smooth system with multiple isolated equilibria that admits a smooth one-to-one linear immersion. These positive and negative examples suggest that the non-existence of continuous one-to-one linear immersions is not solely determined by the presence of multiple isolated equilibria. To provide more accurate guidance on approximating linear immersions, it is imperative to reassess the aforementioned insight and identify the actual factors that determine the existence or non-existence of continuous one-to-one linear immersions.
@@
To address these inquiries, in this work, we study the properties of continuous linear immersions for systems with multiple limit sets, and their implications on algorithms that approximate linear immersions from data. In particular, our contributions include:
  • We introduce a novel class of systems, termed systems with closed basins, including both finite-dimensional linear systems and incrementally stable systems (Section 2).
@@
  • For systems with multiple $\omega$-limit sets, we prove that any continuous immersions into a system with closed basins, under mild conditions, collapses all the $\omega$-limit sets into one and thus can not be one-to-one. We then demonstrate the applicability of our results with multiple examples from the literature (Section 3).
@@
  • For the same class of systems, we show that approximate linear immersions learned with data converge to functions that collapse all the $\omega$-limit sets, as sampling time decreases and sample size increases (Section 4).
@@
A preliminary version of this work was presented at the IFAC World Congress [20], focusing exclusively on one-to-one immersions. In this work, we extend the results in [20] to encompass immersions that are not necessarily one-to-one in Section 3. Additionally, we introduce entirely new results in Sections 4 and 5.
@@
Related work: There is a rich literature on the classes of systems that can be immersed into linear systems, including systems confined in the domain of attraction of a stable equilibrium [10, 17] or a closed orbit [17], poly-flow systems [33], observable nonlinear systems [19], and even control-affine systems (for bilinear immersions) [21]. However, the existence of linear immersions for systems with multiple isolated equilibria remains unresolved. Given that the presence of multiple equilibria is a key feature distinguishing nonlinear systems from linear ones, there has been significant discussion on whether such systems can be immersed into linear systems. It is initially observed in a numerical example from [22] that the approximate linear immersion over a domain that contains two equilibria becomes singular at one of the equilibria. Motivated by this observation, Williams et al. [35] suggest that linear immersions should be approximated within the domain of attraction of each equilibrium to avoid singularities. This suggestion is supported by the work from [24], which studies a one-dimensional system with three equilibria where all the linear immersions can be derived manually. For this specific system, all the derived linear immersions become singular at one of the equilibria. Potentially motivated by these negative examples, Brunton et al. [7] claim that it is impossible to find one-to-one linear immersions for systems with multiple isolated equilibria. However, this claim is disproved by [3], which presents a one-dimensional system with three equilibria that admits a discontinuous one-to-one linear immersion. The paper [3] further conjectures that a linear immersion may exist but become discontinuous at the boundaries of the basins of attraction. This conjecture is again disproved by [2], which constructs a smooth system with two isolated equilibria that admits a smooth one-to-one linear immersion. Contrary to these varying claims, our work rigorously proves that continuous linear immersions cannot be one-to-one when the system has multiple isolated equilibria and satisfies specific conditions. Our results confirm that the negative examples from the literature do not admit continuous one-to-one linear immersions, while the counter example in [2] is the only one not meeting our extra conditions and thus allowing a continuous one-to-one linear immersion. Notably, [16] provides necessary and sufficient conditions for the existence of one-to-one linear immersions for a class of nonlinear systems, which is different than the class of systems with multiple limit sets we consider in this paper. While neither of these two classes is a subset of the other, for any system lying in the intersection of them, both our findings and those results from [16] can infer the nonexistence of continuous one-to-one linear immersions.
@@
Notation: We denote the closure of a set $X$ by $\text{cl}(X)$. The symbols $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, and $\mathbb{R}_{>0}$ denote the real line, the set of non-negative real numbers, and the set of positive real numbers. The symbols $\mathbb{Z}$ and $\mathbb{N}$ denote the set of integers and the set of non-negative integers. A function $\alpha:[0,a)\rightarrow\mathbb{R}_{\geq 0}$, for some $a>0$, belongs to class $\mathcal{K}$ if it is strictly increasing and $\alpha(0)=0$.
@@ 2.1 Problem Statement
Given an initial state $\xi$ of the system in (1), we denote the $\omega$-limit set of $\xi$ in $\mathcal{X}$ by $\omega^{+}(\xi)$, that is the set of all $x\in\mathcal{X}$ satisfying that there exists a sequence $t_{n}\rightarrow\infty$ such that ${\lim_{n\to\infty}\varphi(t_{n},\xi)=x}$ [12].
Next, we introduce the definition of immersion, which generalizes the notion of Koopman eigenfunctions [23].
Definition 1.
@@
where $\psi(t,F(\xi))$ is the solution of $\dot{z}=g(z)$.
If the system $\dot{z}=g(z)$ above is linear, the mapping $F$ is called a linear immersion.
Remark 2.
@@
Remark 3.
Linear immersions are tightly related to Koopman operator theory ([7]): A Koopman eigenfunction $F$ is a (not necessarily continuous) linear immersion that immerses $\dot{x}=f(x)$ in a one-dimensional system $\dot{z}=\lambda z$ for some $\lambda\in\mathbb{R}$. The span of the entries of any linear immersion $F$ is a Koopman invariant subspace.
In the Koopman operator literature, the question of embeddings into finite-dimensional linear systems has been studied in the context of finite-dimensional spaces of observables invariant under the Koopman operator, as discussed in [7, 29]. However, a finite-dimensional Koopman invariant subspace that fully characterizes the Koopman operator may not exist. In our work, we establish the non-existence of immersions of an even more general form, not necessarily arising in this fashion.
Remark 4.
@@
Remark 5.
The term “immersion” is also widely used in the study of differentiable manifolds (see for example [18]), which is unrelated to the immersion of dynamical systems considered in this work.
@@
with state $x=(x_{1},x_{2})$. The system has an unstable equilibrium at the origin and one stable limit cycle on the unit circle, as shown by the phase portrait in Fig. 1. Let $\mathcal{X}=\mathbb{R}^{2}\backslash\{0\}$. Intuitively, one may think a linear immersion does not exist for this system over $\mathcal{X}$ since linear systems cannot have a limit cycle. However, this system does admit a one-to-one linear immersion over $\mathcal{X}$. Let $F:\mathcal{X}\rightarrow\mathbb{R}^{3}$ be
@@
This paper focuses on another well-known topological difference between linear and nonlinear systems, namely that a nonlinear system can have multiple isolated $\omega$-limit sets, but a linear system cannot. We wonder how the properties of linear immersions, such as injectivity, are influenced by the presence of multiple $\omega$-limit sets. Many existing works [22, 7, 24] claim or observe that a continuous one-to-one linear immersion does not exist when the system possesses multiple isolated equilibria, a specific type of $\omega$-limit sets. However, a formal analysis of this phenomenon is missing in the literature. The following discussion explains why it is nontrivial to prove this claim: Suppose that an immersion $F$ that maps a system $\dot{x}=f(x)$ with multiple equilibria to a linear system $\dot{z}=Az$ exists. According to (3), a one-to-one $F$ must map equilibria of $\dot{x}=f(x)$ to equilibria of the linear system $\dot{z}=Az$, and map non-equilibrium points to non-equilibrium points. However, recall that a linear system can only have one isolated equilibrium or a subspace of equilibria. If $\dot{z}=Az$ is the former, then $F$ maps all equilibria of $\dot{x}=f(x)$ to the unique equilibrium of $\dot{z}=Az$ and thus can not be one-to-one. Thus, the main challenge of proving or disproving this claim is to show if it is possible to have a one-to-one $F$ that maps $\dot{x}=f(x)$ to an immersing system $\dot{z}=Az$ with a subspace of equilibria. In this case, the only possibility is that the graph of $F$ intersects with the null space of $A$ at exactly $M$ points, with $M$ the number of equilibria of $\dot{x}=f(x)$, as demonstrated by Fig. 2. The following example shows that this is indeed possible if we allow the immersion to be discontinuous.
Example 2.
Consider a one-dimensional system $\dot{x}=f(x)$ with $M$ isolated equilibria $\{x_{e,i}\}_{i=1}^{M}$, where $x_{e,1}<x_{e,2}<\cdots<x_{e,M}$. Assume that the solution $\varphi(t,\xi)$ is defined for any $\xi$ and $t\in\mathbb{R}$ (and thus has no finite-time blow-up both forward and backward in time). Let $\mathcal{X}=\mathbb{R}$. This system is immersed in the following two-dimensional linear system
@@
$F(x)=$ (8)
$\begin{cases}(x,0)&f(x)=0,\\ (\varphi^{-1}(x,x_{e,1}-1),1)&x<x_{e,1},\\ \left((i+1)\varphi^{-1}\left(x,\bar{x}_{e,i}\right),i+1\right)&x\in(x_{e,i},x_{e,i+1}),\\ ((M+1)\varphi^{-1}(x,x_{e,M}+1),M+1)&x>x_{e,M},\\ \end{cases}$
where $\bar{x}_{e,i}=(x_{e,i}+x_{e,i+1})/2$, and for any $x$ and $\xi\in(x_{e,i},x_{e,i+1})$, the inverse function $\varphi^{-1}(x,\xi)$ is the time instance $t$ such that $\varphi(t,\xi)=x$, which uniquely exists since $x$ and $\xi$ lie on the same trajectory. Intuitively, the equilibria of $\dot{x}=f(x)$ cut the real line into $M+1$ intervals, and the function $F$ in (8) maps trajectories within intervals to different horizontal lines in the lifted space, indexed by the state $v$.
Figure 3: The image of the discontinuous immersion $F$ in (10) is shown by the blue lines and dots. The set of equilibria of the immersed system in (7) is shown by the red line.
To see how this discontinuous immersion $F$ works in practice, we take a concrete example of $\dot{x}=f(x)$ with $M=3$
@@
The equilibria of this system include ±$\pm 1$ and 00. According to (8), a one-to-one discontinuous linear immersion $F$ for this system is
$F(x)=\begin{cases}(x,0)&f(x)=0,\\ \left(\ln\bigg{|}\frac{3x}{-2+2x^{2}}\bigg{|},1\right)&x<-1,\\ \left(2\ln\bigg{|}\frac{3x}{-2+2x^{2}}\bigg{|},2\right)&-1<x<0,\\ \left(3\ln\bigg{|}\frac{3x}{-2+2x^{2}}\bigg{|},3\right)&0<x<1,\\ \left(4\ln\bigg{|}\frac{3x}{-2+2x^{2}}\bigg{|},4\right)&x>1.\\ \end{cases}$ (10)
The image of $F$ is shown in Fig. 3. Let us examine the correctness of $F$ for $x_{0}\in(0,1)$. For any $x_{0}\in(0,1)$, the solution $\varphi(t,x_{0})$ for $\dot{x}=f(x)$ in (9) is
@@
$F(\varphi(t,x_{0}))$ $=\left(3\ln\left(\frac{3x_{0}}{2-2x_{0}^{2}}\right)+3t,3\right).$ (12)
@@
Note that the graph of $F$ in (8) intersects with the subspace ×$\mathbb{R}\times\{0\}$ of equilibria of (7) at precisely $M$ points, thanks to the discontinuity. This phenomenon is also observed in continuous one-to-one linear immersions, exemplified in [2]. This work aims to elucidate the relation between the properties of linear immersions, such as injectivity and continuity, and the occurrence of multiple $\omega$-limit sets, as indicated by these examples and the literature.
Problem: Identify the properties of linear immersions for systems with multiple $\omega$-limit sets.